CRY+VIS Interface Meeting
Date: 2021/8/6 13:00-14:15
Participants: Takahashi, Kimura, Ikeda, Hirata, Sato, Miyo, Ushiba, Aso, Node, Yokozawa, Miyoki
- IX: In finalization. Problem on the cables for payload. Problem of the F2 FR. Installation of the ACCs is scheduled on 8/16-17.
- EX: In finalization. Cable (P7-3) was recovered.
- IY: Height adjustment will be scheduled after vacation.
- EY: Preparation for the payload with the mirror is on going. RM was suspended. Assembly will be finished within this week end. Installation into the cryostat is scheduled on 8/20. Chamber closing is scheduled on 9/6.
- TM height setting: ITMY.
- Commissioning in air: on going in X-arm. Oplev calibration is scheduled on 8/9-11.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Every Friday.
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers:
External review (8/19): improvement in the payload and tower, strategy of height adjustment (Height range), SC plan. Height check of green laser was suggested.
Participants: Takahashi, Aso, Sato, Hirata, Sugiyama, Yano, Fabian, Ishizaki, Tanaka, Terrence
This is generated by python script in /users/Miyo/dropbox/Script/scrapping/main.py.
- Repaired the cable for payload in EX.
- Adjusted the height of GAS filters to the setpoints. The FR for F2 in IX didn't work.
- Prepared for the accelerometers.
The F1 release sensor in EX is malfunctioning. k-log
- In SRM we measured the height of the suspended lower breadboard (LBB):
- Average height: 380.3 mm (nominal 380 m)
- Mass on top: 35.7 kg.
- In SR2 we also measured the height of the LBB at the four suspension points.
- The table is slightly tilted but we didn't change aything becuase the green laser was aligned in O3GK with the steering mirror o the table.
- The height in the area where the mid-size baffle will be is 379 mm.
- The height close to where the steering miror is located is 381 mm.
- Height of the LBB in BS:
- We had to adjust the vertical positions of the hooks becuase the table was 11 mm off.
- After the adjustment the average height became 379.6 mm.
- The ballast mass on the table is 57.5 kg (58 kg rquired by the design).
Cleanliness of SRM: There are about three dirt spots relatively close to the centre of the mirror. In the movie, they can be seen from seconds 10 through 12.
Cleanliness of SR2: There are two spots of dirt at the HR side. They are not so close to the centre. See movie.
- We forgot to check the cleanliness of the BS.
- We moved the OSEM V1 into range. It was not straightforward:
- After the first attempt one OSEM flag was touchig the OSEM body and the oplev was far out of range in pitch.
- We had to move the IM in R, P and Y in order to release the flag and recover the oplev.
- We adjusted the V1 OSEM flag again locking only the RM.
- We measured two IM transfer functions for health assesment.
- We were requested to measure the height of the mirror again in the context of green laser alignment. There was no meaningful change compared to what we measured before.
- We inspected the mirror: there are two stains that are not dust on the HR face. They are not so close to the center but they are not so far away either. See the pictures.
- Type-A team:
- Type-B team: MSB
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
The week from 8/9
- Hirata: 8/9-13
- Yano: 8/9-13
The week from 8/16
- Takahashi: 8/16-20
- Sato: 8/16-20